OpenStudy (anonymous):

For both of the following answer blanks, decide whether the given sequence or series is convergent or divergent. If convergent, enter the limit (for a sequence) or the sum (for a series). If divergent, enter INF if it diverges to infinity, MINF if it diverges to minus infinity, or DIV otherwise. A_n=(9^n)/(-2n+9) The sequence {A_n}?

OpenStudy (anonymous):

It's divergent, I know that for n=1 to infinity

Directrix (directrix):

Looking at these graphs, I'd go with MINF. See what you think: http://www.wolframalpha.com/input/?i=%3D%289%5En%29%2F%28-2n%2B9%29 and http://www.wolframalpha.com/input/?i=Series+%289%5En%29%2F%28-2n%2B9%29

terenzreignz (terenzreignz):

It's divergent all right, and in fact, properly divergent :)

OpenStudy (anonymous):

it says its wrong ;/

Directrix (directrix):

What was given as the correct answer? @monroe17

OpenStudy (anonymous):

It won't show me, I have unlimited attempts until the assignment closes.. I have no idea besides it diverges.

Directrix (directrix):

@monroe17 I think we looked at the series and not the sequence. Hold on. I sent a help request to a Calculus friend. When does the assignment close?

OpenStudy (anonymous):

11:59pm and it's 7:14pm

sam (.sam.):

Wouldn't it be \[\lim_{n \rightarrow -\infty} \frac{9^n}{9-2n}=-\infty\] cause the value gonna get large and in a negative way because of -2n

Directrix (directrix):

@monroe17 From @PhoneAFriend " I wrote out a few terms and the sequence became negatively large very quickly. Check it on my TI-89 and found that by the time we looked at term 100 (or 101, depending on counting technique), the value was already around -1.4 x 10^93. So, the sequence seems to be heading toward negative infinity. Does that seem logical to you?"